1. Consider a disk made of copper, with a specific heat capacity of 0.093 cal/g·C°. How much heat must be removed from a 130 g disk to lower its temperature from 150°C to 24°C? ____ cal

2.A cylinder of gas is heated, with, 870 J of heat is added to the gas. The gas does 160 J of work on its surroundings in this process. What is the change in the internal energy of the gas? ____ J
Answer #1

here,

1)

the specific heat capacity , C = 0.093 Cal /g C

the change in temperature , dT = 24 - 150 degree C

dT = - 126 degree C

mass , m = 130 g

the heat removed , Q = m * C * dT

Q = - 130 * 0.093 * 126 C = - 1523 Cal

the heat removed from the disk is 1523 Cal

2)

the change in internal energy = 870 J - 160 J = 710 J